# Bias in Gini coefficient estimation for gamma mixture populations

### v2

@@ 4 The Gini coefficient estimator is biased for gamma mixtures
$-\sum_{1\leqslant i_{1},\ldots,i_{n-1}\leqslant m}\pi_{i_{1}}\cdots\pi_{i_{n-1}}\,\sum_{j=1}^{m}{n\pi_{j}\alpha_{j}\over\alpha_{i_{1}}+\cdots+\alpha_{i_{n-1}}+\alpha_{j}},$ (26)
because $\mathbb{P}(X=X^{*})=0$.
@@ 5 Illustrative simulation study
$=\left[\sum_{i,j=1}^{m}\widehat{\pi}_{i}\widehat{\pi}_{j}^{*}\,{{}_{2}F_{1}\left(\widehat{\alpha}_{i},-\widehat{\alpha}_{j};\widehat{\alpha}_{i}+1;{1\over 2}\right)\over 2^{\widehat{\alpha}_{i}-1}\widehat{\alpha}_{i}B(\widehat{\alpha}_{i},\widehat{\alpha}_{j}+1)}-1\right]$
@@
$-\left[1-\sum_{1\leqslant i_{1},\ldots,i_{n-1}\leqslant m}\widehat{\pi}_{i_{1}}\cdots\widehat{\pi}_{i_{n-1}}\,\sum_{j=1}^{m}{n\widehat{\pi}_{j}\widehat{\alpha}_{j}\over\widehat{\alpha}_{i_{1}}+\cdots+\widehat{\alpha}_{i_{n-1}}+\widehat{\alpha}_{j}}\right],$ (28)

### v3

@@ 4 The Gini coefficient estimator is biased for gamma mixtures
$-\sum_{1\leqslant i_{1},\ldots,i_{n-1}\leqslant m}\pi_{i_{1}}\cdots\pi_{i_{n-1}}\,\sum_{j=1}^{m}{n\pi_{j}\alpha_{j}\over\alpha_{i_{1}}+\cdots+\alpha_{i_{n-1}}+\alpha_{j}}.$ (26)
@@ 5 Illustrative simulation study
$={1\over n-1}\left[\dfrac{\displaystyle\sum_{1\leqslant i<j\leqslant n}|X_{i}-X_{j}|}{\displaystyle\sum_{i=1}^{n}X_{i}}\right]$
@@
$-\left[1-\sum_{1\leqslant i_{1},\ldots,i_{n-1}\leqslant m}\widehat{\pi}_{i_{1}}\cdots\widehat{\pi}_{i_{n-1}}\,\sum_{j=1}^{m}{n\widehat{\pi}_{j}\widehat{\alpha}_{j}\over\widehat{\alpha}_{i_{1}}+\cdots+\widehat{\alpha}_{i_{n-1}}+\widehat{\alpha}_{j}}\right],\quad n\in\mathbb{N},\,n\geqslant 2,$ (28)